OpenStudy (anonymous):

The length of a Euglena, a small organism and part of the Protist Kingdom, is about 10^-2 centimeters. Express this using positive exponents and evaluate the expression.

OpenStudy (anonymous):

1/10^2

OpenStudy (apoorvk):

\[\large a^-n = \frac{1}{a^n}\] so, 10^(-2) = 1/(10^2) evaluate the length by converting into decimals.

OpenStudy (anonymous):

o.k. thanks I got the same answer just checking it.

OpenStudy (apoorvk):

Very well, should be --> 0.01 cm